# MIND-V: Hierarchical Video Generation for Long-Horizon Robotic Manipulation with RL-based Physical Alignment

### v1

@@ 10 Network Architecture Details
This section provides a detailed specification of the Motor Video Generator (MVG) architecture and its data flow. The MVG is built upon the CogVideoX-5B [cogvideox] model, which features a 3D-VAE and a Diffusion Transformer (DiT) [DiT] backbone. Our primary modification is the conditioning mechanism, which injects guidance from the Behavioral Semantic Bridge (BSB) for precise control. The key specifications and tensor shape transformations are provided in Table 6, with further elaboration below.
@@
The MVG operates within a latent space defined by a 3D Variational Autoencoder (VAE) [vae]. As shown in Table 6, the VAE’s encoder first performs spatiotemporal compression on an input video of shape [B, 3, T, H, W] to produce a compact latent representation of shape [B, 16, T/4, H/8, W/8]. The entire diffusion process is performed within this latent space, with the VAE’s decoder reconstructing the final denoised latent back into pixel space.
@@
The core of the MVG is a 30-layer Diffusion Transformer (DiT) with a hidden dimension of 1920. The latent video is first partitioned into a sequence of 2x2x2 patches, which are then linearly embedded into tokens. This sequence is then augmented with 3D sinusoidal positional encodings and diffusion timestep embeddings before being processed by the Transformer blocks.
@@
The conditioning mechanism integrates BSB guidance directly into the DiT backbone via a Guidance Embedding module. First, the structured information from the BSB is converted into a dense Spatiotemporal Guidance Tensor. As detailed in Table 6, this tensor is then processed by a series of spatiotemporal convolutions—a 3x3 spatial convolution followed by a 1D temporal convolution—to produce a 1920-dimensional guidance signal, $G$. This signal, which retains the same spatiotemporal dimensions as the latent video, is injected into the DiT backbone via additive fusion. The injection occurs within the even-numbered Transformer blocks (0, 2, 4, …), while the odd-numbered blocks remain as standard, unconditioned Transformer blocks. This alternating injection strategy fuses the explicit motion control from the BSB with the video’s internal representations, thereby steering the denoising process to adhere to the planned trajectory.
11 Limitations and Future Work
Despite the promising capabilities of MIND-V in generating physically compliant and semantically consistent manipulation videos, our framework exhibits certain limitations that we actively address.
First, the Staged Visual Future Rollouts mechanism introduces additional computational overhead. The iterative propose-verify-refine cycle requires the generation of multiple candidate videos at each sub-task transition, which increases the inference latency compared to single-pass monolithic models. However, this design choice effectively mitigates the error accumulation inherent in long-horizon tasks. While the per-step inference time is higher, our method achieves high-quality generation with greater overall computational efficiency compared to baselines that often require extensive random sampling or cherry-picking to yield a single successful result. Furthermore, this process operates autonomously and eliminates the need for human intervention during generation.
Second, the framework relies on the accuracy of upstream components within the Semantic Reasoning Hub (SRH), where potential failures in affordance [affordance-r1] localization may propagate to downstream generation. To address this dependency, we incorporate a robust fallback mechanism. In scenarios where the primary visual localizer fails or outputs erroneous results, the VLM directly infers the coordinates of the target object and destination based on semantic context and relative spatial reasoning. Following iterative optimization of these coordinates, the system utilizes them as point prompts for the SAM2 [sam2] to obtain precise segmentation masks.
Future work aims to extend the current 2D video generation framework to 3D representations to better support direct sim-to-real transfer for robotic control.

### v2

@@ 10 Network Architecture Details
This section details the architectural design and data flow of the Motor Video Generator (MVG). Built upon the CogVideoX-5B [cogvideox] foundation, the MVG employs a 3D Variational Autoencoder (VAE) and a Diffusion Transformer (DiT) [DiT] backbone. Our primary modification is the conditioning mechanism, which injects explicit spatiotemporal guidance from the Behavioral Semantic Bridge (BSB) to achieve precise kinematic control. The key module specifications and tensor transformations are summarized in Table 8 and elaborated below.
@@
Latent Space Definition.  The MVG operates within a compressed latent space parameterized by a pre-trained 3D-VAE [vae]. As detailed in Table 8, the VAE encoder performs spatiotemporal compression on an input video of shape $[B,3,T,H,W]$, producing a compact latent representation of shape $B,16,T/4,\allowbreak H/8,W/8]$. The entire forward and reverse diffusion processes are executed within this latent space. Subsequently, the VAE decoder maps the final denoised latent back to the pixel space to render the simulated video.
@@
Diffusion Transformer Backbone.  The core of the MVG simulation engine is a 30-block DiT with a hidden dimension of 1920. The latent video is first partitioned into non-overlapping $2\times 2\times 2$ spatiotemporal patches, which are linearly embedded into a sequence of tokens. This token sequence is augmented with 3D sinusoidal positional encoding and diffusion timestep embeddings prior to being processed by the transformer blocks.
@@
BSB Conditioning Mechanism.  To enforce action adherence, the conditioning mechanism integrates BSB guidance directly into the DiT backbone via a Guidance Embedding module. Initially, the structured semantic information from the BSB is rasterized into a dense Spatiotemporal Guidance Tensor. This tensor is processed by a sequence of spatiotemporal convolutions—specifically, a $3\times 3$ spatial convolution followed by a 1D temporal convolution—yielding a 1920-dimensional feature representation, $G$. This guidance signal, maintaining identical spatiotemporal resolution to the latent video, is injected into the DiT backbone via additive fusion. This injection is performed exclusively within the even-numbered transformer blocks (e.g., blocks 0, 2, 4, …), leaving the odd-numbered blocks unconditioned. This alternating fusion strategy effectively interleaves explicit kinematic control from the BSB with the model’s internal video generation priors, steering the denoising trajectory to strictly adhere to the planned trajectory.